OpenStudy (anonymous):

could someone help me and explain what i need to do ? I'm not sure on how to do this. The amount of cloth used to make four curtains is given by the function A = -4x² + 40x, where x is the width of a curtain in feet and A is the total area of material in square feet. 1) Determine width of the curtain given the total area of the curtain is 96ft². 2) Determine the area of the curtain given a width of 3.5 feet. 3) Graph the function and label the axis of symmetry, vertex, and the roots (solutions). 4) When the curtain is 5ft width, will 60ft² material be enough? .

OpenStudy (mysticfury):

man whoever made this question needs to be a bit more clear on what to do. idk what there asking but id say c is wrong cause it dosent sound right, i say b or d but im not sure

OpenStudy (agreene):

\[A = -4 x^2 + 40x\] 1) A = 96 2) x= 3.5 3 ) graph A(x) 4 ) \(60 = -4(5^2) + 40(5)\) is this true?

OpenStudy (anonymous):

1) Substitute for A and solve for x 2) Substitute for x and solve for A 3) graph y=A(x) 4) Check wether LHS=RHS by substituting both A and x For the first part we have A=96 \[-4x^2+40x=96\] Rearranging a little \[x^2-10x+24=0\] This is a quadratic equation, solve by factoring or using the quadratic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]